Vincent is three times as old as David was two years ago. In three more years, David will be two-thirds as old as Vincent. Neither one is a teenager. How old are Vincent and David now?
Answer by mananth(16946)

let david be x years
2 years ago david was x-2 years
Vincent's age now = 3(x-2)
ages after 3 years
David x+3
Vincent 3(x-2)+3
david age is 2/3 of vincent's age
x+3= 2/3 ( 3 (x-2)+3)
x+3= 2/3 (3x-3)
3(x+3)=2(3x-3)
3x+9=6x-6
3x=15
x=5 David's age
Vincent's age = 3 (x-2)
=3(5-2)
=9 years